OpenStudy (anonymous):

Solve the quadratic equation by completing the square. x^2 + 6x - 5 = 0

OpenStudy (anonymous):

\[x^2+6x=5\] \[(x+3)^2=5+3^2=5+9=14\] \[x+3=\pm\sqrt{14}\] \[x=-3\pm\sqrt{14}\]

OpenStudy (anonymous):

much much easier than the stupid quadratic formula, because the middle term \(6x\) has and even coefficient, half of 6 is 3, and 3 squared is 9, so nice and neat in this case

OpenStudy (amorfide):

x^2 + 6x -5=0 half the coefficient of x put into completing the square format and always subtract the constant inside the bracket squared and you get (x+3)^2 - 9 - 5 = 0 (x+3)^2 - 14=0 add 14 to both sides (x+3)^2=14 square root both sides x+3= +/-root 14 subtract three x=-3+/-root14